# Phase Diagram from Nonlinear Interaction between Superconducting Order and Density: Toward Data-Based Holographic Superconductor

### v1

@@ 1 Introduction
Nonlinear electrodynamics (NED) has been actively applied to various fields of physics. One of the famous examples is the Heisenberg-Euler action, which is physically regarded as a one-loop effective action of the quantum electrodynamics [1]. Also, the representative work of NED is the Born-Infeld theory [2], which turns out to be a low-energy effective action of open strings [3, 4] (See, e.g., [5] for a review). One advantageous property of the Born-Infeld theory is that this can address the divergence of the electric field self-energy of a point particle. This property motivates a lot of research, such as regular black holes [6, 7, 8, 9, 10, 11] and the early universe or accelerating universe in cosmology, e.g., [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Moreover, there are many NED application cases in holography. We will come back to these cases below.
@@
In the holographic model perspective, our investigation is an extension of [69, 50]. In the previous work [50], we gave a particular form of the mass function $M(F^{2})$ and see how this NED interaction deforms the superconducting region in the phase diagram. To interpret this extended NED interaction, we take the viewpoint of [62]. Even though the model of [62] has a different normal state from ours, we accept their interpretation. Our study clarifies how the strongly interacting charge carriers affect the shape of the superconducting dome region. Therefore, our holographic result suggests that this interaction plays a vital role in raising the critical temperature. Thus, such an interaction can be a key to understanding the high-$T_{c}$ superconductor. From the bulk point of view, this is described by a dressed effective mass of the complex scalar by the background electric charges via the nonlinear interaction.
@@ 2 Relaxed Holographic Superconductor with Nonlinear Interaction
Another comment is about the extension of this model. One may consider the higher powers of superconducting order $|\varphi|^{2}$ than the linear one. Such a consideration changes crucially physics in the deep region of the superconducting dome. Our previous work [50] reported that a strange asymptotic Lifshitz scaling shows up in the deep region of the superconducting dome even by a simple form of $M(F^{2})$ up to $(F^{2})^{2}$. Studying the effect from higher powers of $|\varphi|^{2}$ could shed light on clarifying the quantum critical point principle [74]. Nevertheless, the present study excludes this extension for simplicity. Another possible consideration is taking into account $\tilde{F}^{\mu\nu}F_{\mu\nu}$ in the model. Such a consideration plays a crucial role in the magnetic response, such as the Hall angle. We ignore this dependence just for simplicity.
@@
The axion equation of motion is automatically satisfied using (2), so the other equations respect the axion as the relaxation parameter $\kappa$. One can see that this system admits an exact solution known as the Reissner-Nordström (RN)-AdS black hole relaxed by the linear-axion field when $\phi=0$. This solution is dual to the metal (normal) phase, while a hairy solution with nonvanishing $\phi(r)$ describes the superconducting phase.
Now, we demonstrate how to determine the phase boundary of a superconducting dome concerning the critical temperature and charge carrier density. This charge density is regarded as the doping parameter in holographic literature [^3] The reason is that doping changes the charge density of a material. This is a conventional identification in most of the holography literature. In this convention, the optimal doping relies on the superconducting dome shape.. In the presence of a black hole horizon and a scalar hair, the temperature and entropy density of the dual system are defined as the following horizon values:
@@ 3.2 Physics-informed neural network
The hand right side of the eq.(18) is the same as the KAN model, except the basis functions have been replaced from B-splines to Gaussians. The reason we choose Gaussian basis functions is that we do not fix the value of the bare mass $m^{2}$ to be zero. B-splines have a value only within a given interval, whereas Gaussians do not, allowing the $m^{2}$ value to be determined naturally. $\sigma$ is a nonlinear activation function, the positional embedding layer PE is a simple operation, and the neural network operation involves matrix multiplication,
@@
Here, w represents vector $w_{i}$, and $t$ signifies the iteration number in the optimization process, serving as a temporal index that tracks the progression through successive steps. ${\textbf{v}}_{t}$ and ${\textbf{m}}_{t}$ represent the exponentially moving averages of the gradients and the squared gradients, respectively, with ${\textbf{v}}_{t}$ focusing on the direction of the steepest descent. At the same time, ${\textbf{m}}_{t}$ adapts the learning rate based on the historical gradient magnitude. The decay rates $\beta_{1}$ and $\beta_{2}$ are set to 0.99 and 0.999, respectively, with a learning rate $\eta$ of 0.01 and $\epsilon$ at $10^{-8}$ to ensure numerical stability.
@@ 4 Results
Figure 3: Mass function for (a) Gaussian Shape, (b) Double Gaussian shape, (c) Semi-circle shape, (d) Triangle shape phase diagram(Inset figure in each figure).
@@
As a preliminary check of the existence of the hairy solution, including the back reaction, we pick a particular Gaussian shape case and calculate the free energy difference, $\Delta\mathcal{F}=\mathcal{F}_{\text{hairy}}-\mathcal{F}_{RN}$, between its hairy black brane and the RN-AdS black brane. The difference is plotted in Figure 5 (a). The figure’s red line denotes the free energy difference of previous work [50]. The blue line shows the present model, where we set $M(F^{2})=-4\,e^{-(F^{2}+2)^{2}/2}$. In both cases, the free energy of the hairy black brane is always smaller than that of the RN-AdS black brane below the critical temperature; therefore, the hairy solution is energetically preferred at low temperatures. Figure 5 (b) shows the temperature dependence of the dual operator condensation to the scalar field. The condensation behaves $(T-T_{c})^{1/2}$; hence, one can see that the phase transition is a second order.
@@ 5 Conclusion and Future Directions
Then, we tried to find mass functions for two phase-diagram data obtained from experiments. Our main result is shown in Figure 7. The location of phase borderlines is affected by the NED of the broken phase. This implies that the charge-carrier self-interaction, comparable to electron-electron interaction in real materials, has a significant role in determining the critical temperature of the dual field theory. Type I superconductors commonly have low critical temperatures, whereas high-$T_{c}$ superconductors have high critical temperatures. In the view of a holographic method based on our models, it is plausible that superconducting domes accompanied by high critical temperature appear due to the charge-carrier interaction with the superconducting order. From the gravity model perspective, this phenomenon originates from the dressed effective mass of the complex scalar field by the background charge. Based on our AI combined result, one can conjecture that high critical temperatures of superconductors may be related to the electron-electron interaction. This cannot be captured in the band theory framework, which relies on the single-particle periodic potential. It would be valuable to study this physics in both theories of holography and condensed matter.

### v2

@@ 1 Introduction
Nonlinear electrodynamics (NED) has been actively applied to various fields of physics. One of the famous examples is the Heisenberg-Euler action, which is physically regarded as a one-loop effective action of the quantum electrodynamics [1]. Also, the representative work of NED is the Born-Infeld theory [2], which turns out to be a low-energy effective action of open strings [3, 4] (See, e.g., [5] for a review). One advantageous property of the Born-Infeld theory is that this can address the divergence of the electric field self-energy of a point particle. This property motivates a lot of research, such as regular black holes [6, 7, 8, 9, 10, 11] and the early universe or accelerating universe in cosmology, e.g., [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Moreover, there are many NED applications in holography. We will come back to such works below.
@@
In the holographic model perspective, our investigation is an extension of [69, 50]. In the previous work [50], we gave a particular form of the mass function $M(F^{2})$ and see how this NED interaction deforms the superconducting region in the phase diagram. To interpret this extended NED interaction, we take the viewpoint of [62]. Even though the model of [62] has a different normal state from ours, we accept their interpretation. Our study clarifies how the strongly interacting charge carriers affect the shape of the superconducting dome region. Therefore, our holographic result suggests that this interaction plays a vital role in raising the critical temperature. Thus, such an interaction can be a key to understanding the high-$T_{c}$ superconductor. From the bulk point of view, this is described by a dressed effective mass of the complex scalar by the background electric charges via the nonlinear interaction. We will discuss this issue later.
@@ 2 Relaxed Holographic Superconductor with Nonlinear Interaction
Another comment is about the extension of this model. One may consider the higher powers of superconducting order $|\varphi|^{2}$. Such a consideration changes crucially physics in the deep region of the superconducting dome. Our previous work [50] reported that a strange asymptotic Lifshitz scaling shows up in the deep region of the superconducting dome even by a simple form of $M(F^{2})$ up to $(F^{2})^{2}$. Studying the effect from higher powers of $|\varphi|^{2}$ could shed light on clarifying the quantum critical point principle [74]. Nevertheless, the present study excludes this extension for simplicity. Another possible consideration is taking into account $\tilde{F}^{\mu\nu}F_{\mu\nu}$ in the model. Such a consideration plays a crucial role in the magnetic response, such as the Hall angle. We ignore this dependence just for simplicity.
@@
The axion equation of motion is automatically satisfied using (2), so the other equations respect the axion as the relaxation parameter $\kappa$. One can see that this system admits an exact solution known as the Reissner-Nordström (RN)-AdS black brane relaxed by the linear-axion field when $\phi=0$. This solution is dual to the metal (normal) phase, while a hairy solution with nonvanishing $\phi(r)$ describes the superconducting phase.
Now, we demonstrate how to determine the phase boundary of a superconducting dome concerning the critical temperature and charge carrier density. In the gauge/gravity duality, the conserved charge density is obtained from the bulk Maxwell equations. This conserved charge plays a role in the charge carrier density from the holographic DC electric conductivity. In the high $T_{C}$ cuprate experiment, the doping parameter is used as a control parameter. The role of doping is to add an electron or hole by replacing an atom in the material. Therefore, changing the doping parameter implies changing charge carrier density, and both are proportional. In this sense, we regard charge carrier density $Q$ as a doping parameter [^3] It is not a unique way to define doping parameters holographically. In previous work[53], for example, they introduce another $U(1)$ gauge field, which corresponds to the doping parameter. In this case, they can introduce charge carriers and doping parameters individually. In other work[54], they defined doping parameters as a ratio of two conserved charges; one is for the charge carriers, and the other is for impurities. .
In the case of the high $T_{C}$ cuprate, the zero doping region is covered by an anti-ferromagnetic insulating phase. This phase is insulating; hence, the charge carrier density should be zero. This insulating phase still exists at the finite doping region; therefore, the charge carrier density and the doping parameter are not exactly proportional. To match the experimental parameter to the model parameter, we have to take care of all dimensionful parameters and carefully find parameter mapping.
In the presence of a black hole horizon and a scalar hair, the temperature and entropy density of the dual system are defined as the following horizon values:
@@ 3.2 Physics-informed neural network
The right-hand side of the eq.(18) is the same as the KAN model, except the basis functions have been replaced from B-splines to Gaussians. The reason we choose Gaussian basis functions is that we do not fix the value of the bare mass $m^{2}$ to be zero. B-splines have a value only within a given interval, whereas Gaussians do not, allowing the $m^{2}$ value to be determined naturally. $\sigma$ is a nonlinear activation function, the positional embedding layer PE is a simple operation, and the neural network operation involves matrix multiplication,
@@
Here, w represents vector $w_{i}$, and $t$ signifies the iteration number in the optimization process, serving as a temporal index that tracks the progression through successive steps. ${\textbf{v}}_{t}$ and ${\textbf{m}}_{t}$ represent the exponentially moving averages of the gradients and the squared gradients, respectively, with ${\textbf{v}}_{t}$ focusing on the direction of the steepest descent. At the same time, ${\textbf{m}}_{t}$ adapts the learning rate based on the historical gradient magnitude. The decay rates $\beta_{1}$ and $\beta_{2}$ are set to 0.99 and 0.999, respectively, with a learning rate $\eta$ of 0.01 and $\epsilon$ at $10^{-8}$ to ensure numerical stability. We summarize the structure of our neural network in Figure 1.
@@ 4 Results
Figure 3: Mass functions for (a) Gaussian Shape, (b) Double Gaussian shape, (c) Semi-circle shape, (d) Triangle shape phase diagram (Inset figure in each figure).
@@
As a preliminary check of the existence of the hairy solution, including the back reaction, we pick a particular Gaussian shape case and calculate the free energy difference, $\Delta\mathcal{F}=\mathcal{F}_{\text{hairy}}-\mathcal{F}_{RN}$, between its hairy black brane and the RN-AdS black brane. The difference is plotted in Figure 5 (a). The figure’s red line denotes the free energy difference of previous work [50]. The blue line shows the present model, where we set $M(F^{2})=-4\,e^{-(F^{2}+2)^{2}/2}$. In both cases, the free energy of the hairy black brane is always smaller than that of the RN-AdS black brane below the critical temperature; therefore, the hairy solution is energetically preferred at low temperatures. Figure 5 (b) shows the temperature dependence of the dual operator condensation to the scalar field. The condensation behaves $(T-T_{c})^{1/2}$; hence, one can see that the phase transition is a second-order phase transition.
@@
Comparing the actual data (Figure 6) and our results (Figure 7), one can notice that the numbers look different. However, since the momentum relaxation parameter $\kappa$ values are different for each material, and the doping parameter and the charge carrier density $Q$ are proportional, but the proportionality constant is also different for each material, the discrepancy in the numbers on the axes can be eliminated by choosing these parameters at any time. We express the numbers on the axes as dimensionless quantities to honestly show this material-dependent ambiguity.
@@
The effective mass obtained by training fits experimental data very well in Figure 7, but the exact form of the effective mass is very complicated. One can notice that the phase diagram consists of two peaks with different positions and widths. We propose a mass function as a combination of two Gaussian functions as
$M(F^{2})=A_{1}\,e^{\frac{-(F^{2}-B_{1})^{2}}{{\Delta_{1}}^{2}}}+A_{2}\,e^{\frac{-(F^{2}-B_{2})^{2}}{{\Delta_{2}}^{2}}},$ (23)
[…]
@@ 5 Conclusion and Future Directions
Then, we tried to find mass functions for two phase-diagram data obtained from experiments. Our main result is shown in Figure 7. The location of phase borderlines is affected by the NED of the broken phase. This implies that the charge-carrier self-interaction, comparable to electron-electron interaction in real materials, has a significant role in determining the critical temperature of the dual field theory. Type I superconductors commonly have low critical temperatures, whereas high-$T_{c}$ superconductors have high critical temperatures. In the view of a holographic method based on our models, it is plausible that superconducting domes accompanied by high critical temperature appear due to the charge-carrier interaction with the superconducting order.
On the other hand, from the gravity model perspective, this phenomenon originates from the background charge. Due to the nonlinear interaction between the complex scalar field and the Maxwell field, the scalar field feels its mass with the dressing effect from the background charge. This effective mass becomes bare mass near the asymptotic boundary. However, the dressing mass effect shows up more and more as the scalar field approaches the horizon. Therefore, the scalar field becomes more unstable near the horizon. This helps the scalar hair form. As a result, the critical temperature becomes higher.
Based on our AI combined result, one can conjecture that high critical temperatures of superconductors may be related to the electron-electron interaction dual to the complex scalar dressing mass effect. This cannot be captured in the band theory framework, which relies on the single-particle periodic potential. It would be valuable to study this physics in both theories of holography and condensed matter.
@@
In this work, we focus on the mass function, which gives a similar shape to the phase diagram with experimental data. we set Newton’s constant $G$ and the AdS radius $L$ to one. And we also rescale every coordinate to the black hole horizon radius. To match our result to the experimental data quantitatively, we have to consider all dimensionful parameters. This work is a starting point for applying machine learning techniques to the gauge/gravity duality and whether the process is working well. The background geometry for normal phase does not give a strange metallic behavior. Therefore, we postpone the parameter matching process as a future work.